# InsightBench: Evaluating Business Analytics Agents Through Multi-Step Insight Generation

### v1

@@ 6 Conclusion
We have introduced InsightBench, a benchmark with 31 diverse datasets that evaluates agents on their ability to perform end-to-end data analytics, including suggesting questions, interpreting answers, and summarizing insights. We ensured each dataset has clear goals and meaningful analysis as ground-truth to evaluate agents reliably. Using LLaMA-3-Eval, an open-source evaluator, we assessed how well the agents extracted insights compare to the ground-truth. We showed how our proposed agent, AgentPoirot, can perform data analytics from high-level goals using both open-source and closed-source models and showed that it outperforms existing approaches like Pandas Agent. We believe InsightBench will significantly drive advancements in data analytics. For future work, we look into expanding this benchmark to include more categories of data such as healthcare data, social media trends, environmental data, e-commerce analytics, and educational statistics.
@@ Appendix A Limitations and Potential Societal Impact
Diversity and Scope of Data Simulation.    While InsightBench covers a spectrum of business areas and themes, the datasets are synthesized based on patterns observed in widely-used ServiceNow. Consequently, the synthetic "flags" are designed to mimic real-world anomalies and trends but may not capture genuine enterprise data’s full complexity or unexpected behavior.
@@ Appendix B Appendix: Overview
B.1 Accessibility and Maintenance
Table 3: Following the dataset track guidelines, we share the following availability of InsightBench dataset
Artifact Link Licence
Home Page https://insightbench.github.io/ CC BY-NC 4.0
Code Repository https://github.com/ServiceNow/insight-bench MIT License
Datasets https://huggingface.co/datasets/ServiceNow/insight_bench CC BY-NC 4.0
@@ B.1 Accessibility and Maintenance
B.2 Example Protocol for Data Generation and Insight Planting
@@
    Schema Exporting and Standardization: Initially, the schema for the incidents table is exported and standardized by inspecting a demo instance. Fields such as "number", "opened_at", and "closed_at", among others, are defined.
@@
Table 4: Overview of InsightBench Datasets and Problems. This table enumerates all 31 datasets included in the benchmark, detailing their respective topics, difficulty levels, and titles. This enlists diverse scenarios that InsightBench covers to evaluate the abilities of data analytics agents.
@@
Goal Management in a Department Analysis 31 4 Goal Management
@@
B.3 Data Tables across themes
@@
B.4 List of Datasets and Problems
@@
B.5 Quality-check Interface
To ensure the high standards of data quality and relevance, InsightBenchhas implemented a comprehensive quality-check process involving volunteer contributors. This process is facilitated through a specifically developed gradio interface, which guides the expert reviewers through a structured review of the ground-truth analysis notebooks included in the benchmark (Figure˜5). The evaluation is divided into three main sections:
@@ B.5 Quality-check Interface
B.6 Details of the proposed AgentPoirot
@@ Appendix C Ablation Results
(a) Effect of varying sampling temperature (performance reported across the 31 datasets in InsightBench)
@@ Appendix D Prompt Design
For the json file must have a "name", "description", and "value" field that describes the data.
@@
gpt_fields = ["short_description", "assignment_group"]
@@
"<field>"
+ f"<name>{f}</name><type>{col_info[f][’type’]}</type>"
f"<choices>{col_info[f][’choices’]}</choices>"
if "choices" in col_info[f]
else ""
+ "</field>"
@@
{{"short_description": "My laptop is broken"}}
@@
"role": "system",
"content": "Youre going to help me generate simulated data that looks like it really came from ServiceNow Glide tables.",
{"role": "user", "content": prompt},
@@
if "choices" in col_info[f]:
assert incident[f] in col_info[f]["choices"]
@@
print(f"… {len(valid_data)} / {num_samples}")
@@
print("… invalid value, trying again")

### v2

@@ 6 Conclusion
We have introduced InsightBench, a benchmark with 100 diverse datasets that evaluates agents on their ability to perform end-to-end data analytics, including suggesting questions, interpreting answers, and summarizing insights. We ensured each dataset has clear goals and meaningful analysis as ground-truth to evaluate agents reliably. Using LLaMA-3-Eval, an open-source evaluator, we assessed how well the agents extracted insights compare to the ground-truth. We showed how our proposed agent, AgentPoirot, can perform data analytics from high-level goals using both open-source and closed-source models and showed that it outperforms existing approaches like Pandas Agent. We believe InsightBench will significantly drive advancements in data analytics. For future work, we look into expanding this benchmark to include more categories of data such as healthcare data, social media trends, environmental data, e-commerce analytics, and educational statistics.
@@ 7 Limitations
8 Reproducibility Statement
To facilitate the reproducibility of our results, we have taken the following measures:
1. Code Availability: All our code, including data used for model implementation, training, and evaluation, is available in the supplementary material.
2. Computing Infrastructure: All LLaMA-3 experiments were conducted on 2 x 80G A100 GPUs.
4. Ablations: A complete list of hyperparameters used in our experiments is provided in Appendix C of the paper. Additionally, all prompts used in this work are included in Appendix D.
5. Random Seeds: To ensure reproducibility, we have run our experiments for 5 random seeds.
6. Evaluation Metrics: We provide a detailed description of all evaluation metrics used in Section 2.2 of the paper. Implementation of these metrics is included in our codebase.
7. Experimental Procedures: Section 3 of the paper includes a comprehensive description of our experimental procedures.
8. Dependencies: A complete list of software dependencies, is provided in the ‘requirements.txt‘ file of the supplementary material.
By specifying these details, we aim to ensure that our work can be readily reproduced and built upon by the research community.
@@ Appendix A Limitations and Potential Societal Impact
Diversity and Scope of Data Simulation.    While InsightBench covers a spectrum of business areas and themes, the datasets are synthesized based on patterns observed in widely-used ServiceNow. Consequently, the synthetic ”flags” are designed to mimic real-world anomalies and trends but may not capture genuine enterprise data’s full complexity or unexpected behavior.
@@ Appendix B Appendix: Overview
B.1 Example Protocol for Data Generation and Insight Planting
@@
    Schema Exporting and Standardization: Initially, the schema for the incidents table is exported and standardized by inspecting a demo instance. Fields such as ”number”, ”opened_at”, and ”closed_at”, among others, are defined.
@@
Table 3: Overview of InsightBench Datasets and Problems. This table enumerates a subset of the first 30 datasets included in the benchmark, detailing their respective topics, difficulty levels, and titles. This enlists diverse scenarios that InsightBench covers to evaluate the abilities of data analytics agents.
@@
B.2 Data Tables across themes
@@
B.3 List of Datasets and Problems
@@
B.4 Quality-check Interface
To ensure the high standards of data quality and relevance, InsightBenchhas implemented a comprehensive quality-check process involving volunteer contributors. This process is facilitated through a specifically developed gradio interface, which guides the expert reviewers through a structured review of the ground-truth analysis notebooks included in the benchmark (Figure 5). The evaluation is divided into three main sections:
@@ B.4 Quality-check Interface
B.5 Details of the proposed AgentPoirot
@@ Appendix C Ablation Results
(a) Effect of varying sampling temperature (performance reported across the 100 datasets in InsightBench)
@@ Appendix D Prompt Design
For the json file must have a name”, description”, and value field that describes the data.
@@
gpt_fields = [”short_description”, assignment_group”]
@@
”<field>”
+ f”<name>{f}</name><type>{col_info[f][’type’]}</type>”
f”<choices>{col_info[f][’choices’]}</choices>”
if choices in col_info[f]
else ””
+ ”</field>”
@@
{{”short_description”: My laptop is broken”}}
@@
role”: system”,
content”: Youre going to help me generate simulated data that looks like it really came from ServiceNow Glide tables.”,
{”role”: user”, content”: prompt},
@@
if choices in col_info[f]:
assert incident[f] in col_info[f][”choices”]
@@
print(f”… {len(valid_data)} / {num_samples}”)
@@
print(”… invalid value, trying again”)